# URLLC for 6G Enabled Industry 5.0: A Taxonomy of Architectures, Cross Layer Techniques, and Time Critical Applications

### v1

@@ IV-A Reliability-Latency-Rate Tradeoff and QoS in URLLC Systems
In [62], Noor-A-Rahim et al. conceptualized the integration of IRSs into A.5 to enable URLLC. The authors presented multiple IRS-assisted use cases, including blockage mitigation, mmWave/THz communications, wireless energy transfer, sensing and localization, and MEC T.1.
@@
The study addresses two major challenges: a) the trade-off between latency and energy efficiency in power constrained UAV platforms (C.1, P.3); and b) maintaining ultra reliability during UAV trajectory changes (C.4) and remote AIoT device operations (P.2). To overcome these issues, the authors developed a low complexity optimization framework that jointly manages resource allocation and UAV trajectory planning. The optimization is based on a successive convex approximation algorithm combined with a random restart method to escape local minima and improve convergence. The model aims to minimize the weighted sum of energy consumption across UAVs and ground AIoT devices while satisfying URLLC level constraints, such as sub-millisecond latency and ultra low packet error rates. Specifically, smaller blocklengths and tighter error tolerances are incorporated into the QoS formulation to ensure deterministic transmission bounds. The proposed work has been shown to reduces total energy consumption P.3 compared to benchmark fixed trajectory and static resource allocation strategies. This work enhances URLLC support in dynamic edge environments by ensuring end-to-end reliability (P.2), achieving latency constrained energy savings (P.3). These capabilities make the proposed framework suitable for real-time Industry 5.0 deployments involving mobile AIoT networks and mission critical edge applications.
@@
In [63], Imtiaz et al. proposed a DT enabled MEC framework aimed at improving computational energy efficiency in latency sensitive consumer Industrial IoT (IIoT) applications for Industry 5.0 smart manufacturing environments (A.5) to balance latency vs. reliability trade-offs C.1. The system as shown in Figure 7 introduces predictive digital twin agents that continuously synchronize with physical edge computing nodes, enabling proactive decision making for task offloading. Unlike traditional reactive MEC systems, the proposed framework predicts resource bottlenecks in advance and dynamically schedules tasks to minimize latency and energy consumption simultaneously (T.5, T.1).
@@
The key optimization challenge is modeled as a mixed integer non-linear programming problem that jointly minimizes energy consumption and task latency while satisfying strict URLLC constraints. To achieve computational tractability, the authors apply piecewise linear approximation to relax non-linear constraints. Task assignment decisions are made based on predicted CPU availability, wireless channel quality, and residual energy of nodes, allowing latency aware offloading while ensuring reliability. This predictive scheduling framework enhances both the responsiveness and sustainability of real-time IIoT applications by allocating edge resources more efficiently P.3). The proposed work has been shown to improve computational energy efficiency compared to traditional edge only models and reduces the service outage probability under strict latency and reliability requirements. These improvements contribute to end-to-end latency reduction (P.1), reliability improvement (P.2), and resource optimization (P.3) which are, key performance enhancements required in URLLC systems. The proposed DT assisted optimization approach provides energy efficient foundation for deploying URLLC enabled smart manufacturing services in Industry 5.0.
@@
Li et al. [55] proposed a low latency terminal mutual access architecture based on 5G LAN cross-domain networking designed to support real-time communication demands in dynamic industrial scenarios. The architecture addresses the shortcomings of traditional communication frameworks that route data through central servers, introducing latency and single points of failure. Instead, their solution enables D2D communication using 5G LAN capabilities, which is important for latency sensitive applications such as remote surgeries (A.3), and supporting the evolution toward next-generation mobile networks A.6
The proposed architecture addresses two main challenges: a) enabling real-time, ultra reliable communication across multiple 5G LAN domains managed by different operators (C.2); and b) maintaining privacy and security while bypassing centralized cores (C.3). The architecture integrates MEC T.1 to host computation intensive tasks near end users, network slicing T.2 to logically separate services, and NFV and SFC T.3 to flexibly deploy and interconnect virtual network functions across domains. To achieve this, the authors use a distributed mesh topology with user plane function based routing (i.e., decentralized traffic anchoring at the edge to enable direct data forwarding between terminals without passing through the core network) supported by secure access service edge and edge offloading (T.1). These mechanisms enable traffic to bypass base stations and core networks (C.4), thereby achieving latency isolation and privacy preserving communication under a multi-domain setup (C.3).
@@ IV-F Cross-Layer Networking Solutions
Gao et al. [68] addressed the challenge of delivering URLLC compliant services for latency sensitive metaverse applications such as real-time holographic communication (A.4). The primary difficulty lies in simultaneously satisfying ultra-low latency, high reliability, and immersive QoE requirements under limited wireless and computing resources (C.1, C.4). To address this, the authors proposed a multi-objective optimization framework for joint allocation of communication, computation, and storage resources in edge assisted networks (T.1). The system model considers a heterogeneous user base with varying latency budgets, traffic demands, and task data sizes. Reliability is enforced using probabilistic bounds on decoding errors  (P.2, P.1). The resulting non-convex, integer constrained problem is solved using a decomposition based evolutionary algorithm to approximate the Pareto front, enabling designers to select operating points that best balance URLLC constraints and QoE. Simulation results demonstrate improved delay guarantees (P.1) and higher adaptability (P.3).
@@
In dynamic next-generation wireless networks (A.6), accurate channel prediction is fundamental for optimizing communication and ensuring low latency, reliability and security of data transmission. AI, Including DL and RL, offer great advantages for URLLC use cases. For instance, Li et al. [54] proposed a DRL based joint scheduler that dynamically balances QoS trade-offs between eMBB and URLLC A.6. Unlike static or risk sensitive models, the authors formulated a multi-objective optimization problem using reliability (P.2) and service satisfaction level as key metrics. Reliability is quantified through the probability of successful URLLC packet decoding under finite blocklength constraints, where decoding error probabilities exceeding a target threshold (e.g., $10^{-5}$) adds penalties in the reward. The environment is modeled such that the state includes channel quality indicators and queue lengths, the action consists of selecting resource block assignments and puncturing decisions for URLLC packets, and the reward function jointly optimizes service satisfaction level fairness for eMBB and packet error rate constraints for URLLC traffic. They used Deep Deterministic Policy Gradient (DDPG) with prioritized experience replay to train the scheduler on real-time traffic variations. The proposed approach has been shown to increase service satisfaction level (P.3) and maintain reliability requirements for URLLC transmissions (P.2). By using DRL (T.4), the scheduler learns to reallocate resources under stochastic traffic demands and fluctuating wireless channel conditions (C.4), thereby enabling latency aware, priority preserving access in mixed traffic scenarios. This supports the goals of URLLC by ensuring intelligent resource orchestration, reduced queuing delays, and improved service continuity under hard latency and reliability constraints (P.1).
proposed a DRL based joint scheduler designed to dynamically balance QoS tradeoffs between eMBB and URLLC traffic in next-generation wireless networks A.6. Their primary motivation was to address the challenge posed by the 5G new radio’s puncturing approach, where URLLC traffic is prioritized by overlapping ongoing eMBB transmissions, leading to significant rate loss and degraded QoS for eMBB users (C.4). The authors acknowledge that the cost of eMBB QoS loss could sometimes higher than the gains for URLLC users, leading to the need for a long-term QoS tradeoff by observing environmental variations (P.1). The proposed idea is to jointly optimize bandwidth allocation and the overlapping (i.e., puncturing) positions of URLLC users’ traffic. They formulated this as a multi-objective optimization problem, which was then transformed into a single-objective optimization problem using the objective product method to find Pareto optimal solutions. The decision making problem is modeled as a MDP. The authors used a Deep Deterministic Policy Gradient (DDPG) algorithm with prioritized sampling to handle continuous state and action spaces and improve learning stability.
The DDPG agent at each time slot observes, the channel conditions represented by the signal-to-noise ratio (SNR) for both eMBB and URLLC users, traffic requirements, constant bit rate requirements for eMBB users, and the size of arriving URLLC packets for each user in each time slot. Then the agent takes a joint action that determines, Resource Block (RB) allocation for eMBB users, RB allocation for URLLC users, and puncturing decisions, which are the fraction of RBs that an eMBB user contributes (i.e., is punctured) to URLLC users in time slot. This quantifies the eMBB sacrifice for URLLC. The reward represents: a) the user centric trad-off, which aims to maximize the sum of logarithmic service satisfaction levels. This service satisfaction levels is a user centric metric, indicating if individual rate requirements (e.g., for eMBB) or transmission demands (e.g., for URLLC) are met for both eMBB and URLLC users; b) system wide tradeoff, which seeks to maximize the product of the system wide reliability (P.2) metrics for eMBB and URLLC.
The actor component uses a Deep Neural Network (DNN) that acts as the policy function, mapping observed states to specific actions (i.e., RB allocations and puncturing decisions). The goal of the agent is to learn a policy that maximizes the discounted long-term reward. The critic component also uses a DNN to maintain a state action value function, which evaluates the long-term return of taking a particular action in a given state under the actor’s policy ([36]). This evaluation guides the actor’s learning process. A key enhancement in this work is the use of prioritized sampling in the replay buffer. Instead of uniformly sampling past experiences, transitions with higher loss values (i.e., indicating that the model made a larger error) are sampled more frequently. This helps the agent learn more effectively from significant experiences and accelerates convergence. The proposed work has been shown to achieve better system wide tradeoff scheduling compared to user centric tradeoff methods in terms of URLLC reliability. This study contributes to enabling URLLC by providing an AI enhanced scheduling framework (T.4) for next-generation mobile networks (A.6) that addresses the design challenges of mobility (C.4) and the fundamental latency vs. reliability tradeoffs (P.1) by improving end-to-end latency reduction (P.1) and reliability (P.2).
@@
Similar to [54] and [70], She et al. [34] proposed a multi-level DL architecture that addresses the challenge of meeting URLLC performance requirements (C.4) in non-stationary 6G networks (A.6). Figure 11 shows the proposed framework, which integrates device level intelligence for mobility prediction, edge level (T.1) intelligence for scheduler design, and cloud level intelligence for user association. At the device level, they use fully connected DNNs trained on tactile mobility traces to predict user movements, achieving a prediction error probability below $10^{-5}$, thereby enhancing reliability guarantees (P.2). At the cell level, a DRL based actor critic scheduler is developed and evaluated in a live LTE testbed. The scheduler is trained using two DNNs (i.e., actor and critic), each with multiple hidden layers, to minimize packet loss and ensure delay and jitter bounds. They demonstrated real-time performance under strict constraints using pre-training in a digital twin, then fine tuning in real-world deployments (T.5, P.1). At the network level, the system uses DNN based user association, trained with large scale channel gains and packet arrival rates, and compared against optimal and baseline models. To address learning in dynamic environments, the authors employed deep transfer learning to adapt pre-trained models with limited new data (T.4), and applied hierarchical federated learning to enable distributed training while preserving privacy (C.3) and reducing communication overhead (C.4). The proposed approach has been shown to improve resource utilization (P.3 and energy efficient user association across varying service types and user distributions (P.4, P.5). By extending the learning from device to edge, and cloud layers, this study supports the main principles of URLLC by enabling ultra low latency task scheduling, accurate user mobility tracking, and proactive resource orchestration in highly dynamic environments which ensures reliable performance across heterogeneous traffic scenarios (C.1).
@@
Zhou et al. [28] addressed the challenge of supporting URLLC for smart healthcare services in Internet of Health Things (IoHT) networks (A.3, A.6). Ensuring ultra low latency and high reliability for tasks such as remote health monitoring and emergency diagnosis is challenging due to device heterogeneity, dynamic channel conditions, constrained computation (C.1, C.4), and the need for privacy preserving solutions (C.3). The authors proposed a learning based task offloading approach that uses federated DRL model to optimize computation offloading decisions under URLLC constraints. The model is based on the exponential weight algorithm for exploration and exploitation to optimize computation offloading decisions under URLLC constraints. The proposed approach adopts a hybrid architecture where IoHT devices decide whether to locally execute or offload tasks to edge servers. A federated DQN is used to train local agents collaboratively without sharing raw data, preserving patient privacy (C.3). Each agent observes a state composed of local queue length, channel quality, and device energy. The action includes binary offloading decisions and edge server selections. The reward function includes task success rate and delay violations in order to enforce URLLC compliance. Agents update local policies using prioritized experience replay and synchronize with a global model to adapt to environmental changes (T.4. The proposed approach has been shown to reduce average task latency (P.1) and improve offloading success rate (P.2) compared to traditional approaches. This enables URLLC support for delay sensitive healthcare applications such as real-time ECG monitoring and AR assisted surgery. By integrating privacy aware learning, energy-efficient offloading, and delay sensitive scheduling, the work contributes to reliable and low latency healthcare delivery in dense IoHT environments.
@@
Kokkinis et al. [49] addressed the challenge of synchronizing video and haptic streams under URLLC constraints in tactile Internet environments, which falls under the application domain of Remote Control and Haptics (A.4) [9, 14, 15]. A key issue they focused on was meeting distinct service level requirements for haptic and video traffic, such as sub-10 ms latency for haptic control loops and desynchronization thresholds below 50 ms for user perception, while maintaining spectral efficiency and robustness amidst fluctuating wireless conditions. These challenges inherently involve latency vs. reliability trade-offs (C.1) and mobility management (C.4) [9]. The authors proposed a soft actor-critic based Deep Reinforcement Learning (DRL) framework, leveraging AI/ML for URLLC (T.4), to optimize radio resource allocation between haptic (URLLC slice) and video (eMBB slice) modalities. The DRL environment models state by capturing average latency, packet loss, data rate, spectrum efficiency, and buffer occupancy per slice, with actions corresponding to discrete allocations of resource blocks. Their reward function penalizes violations of service-specific constraints, including slice-wise data rate (P.3), packet loss (P.2), and inter-modal synchronization (P.1). The system reduces computational complexity in edge deployments by performing inter-slice scheduling every 50 ms instead of 1 ms. The proposed framework has been shown to achieve a higher satisfaction rate (P.4) under high user density (P.5) and spectrum efficiency fluctuations compared to traditional approaches. By modeling heterogeneous service demands and optimizing resource allocation via DRL, their approach enables stable, low latency, and synchronized teleoperation services, **contributing to the evolving capabilities of next-generation mobile networks** (A.6).
kokkinis et al. [49] tackled the unique challenge of synchronizing video and haptic streams under URLLC in tactile internet environments (A.4). One of the key issues is meeting distinct service level requirements of haptic and video traffic such as sub-10 ms latency for haptic control loops and desynchronization thresholds below 50 ms for user perception, as well as maintaining spectral efficiency and robustness under fluctuating wireless conditions (C.1, C.4).
The authors proposed a soft actor-critic based DRL framework to optimize radio resource allocation between haptic (i.e., URLLC slice) and video (i.e., eMBB slice) modalities. The environment is modeled as a MDP where the state captures average latency, packet loss, data rate, spectrum efficiency, and buffer occupancy per slice. The action space corresponds to discrete allocation ratios of resource blocks per slice. The reward function is designed to penalize violations of service specific constraints, including slice wise data rate (P.3), packet loss (P.2), and inter-modal synchronization (P.1). The actor-critic agent learns slicing strategies through entropy regularized policy updates, with network outputs mapped to resource block allocations. The system reduces complexity by performing inter-slice scheduling every 50 ms instead of 1 ms, lowering computational burden in edge deployments. The proposed work has been shown to achieve higher satisfaction rate (P.4 compared to traditional approaches under high user density (P.5) and spectrum efficiency fluctuations. By explicitly modeling heterogeneous service demands and jointly optimizing resource allocation using DRL, the proposed approach supports stable, low latency and synchronized teleoperation services. This contributes to enabling URLLC in remote control and haptic scenarios by ensuring adaptive, learning based slicing across time varying radio environments (A.6.
@@
The above studies demonstrate that AI and ML techniques are increasingly central to enabling URLLC in 6G enabled Industry 5.0 applications. From optimizing dynamic slicing and uplink scheduling to managing delay constrained offloading and haptic video synchronization, AI agents provide scalable and adaptive control policies under stochastic traffic and wireless dynamics (T.4). These learning based approaches: a) improve latency reliability trade-offs (C.1); and b) supports intelligent orchestration across mobility intensive (C.4) and heterogeneous service environments. This is done by bridging real-time perception, decision-making, and control, in order to contribute to fulfilling the end-to-end URLLC objectives of latency reduction (P.1), reliability improvement (P.2), and resource efficiency (P.3, P.4), which in turn makes it one of the main enablers of intelligent and responsive Industry 5.0 systems.

### v2

@@ IV-A Reliability-Latency-Rate Tradeoff and QoS in URLLC Systems
In [62], Noor-A-Rahim et al. conceptualized the integration of IRSs into (A.5) to enable URLLC. The authors presented multiple IRS-assisted use cases, including blockage mitigation, mmWave/THz communications, wireless energy transfer, sensing and localization, and MEC (T.1).
@@
The study addresses two major challenges: a) the trade-off between latency and energy efficiency in power constrained UAV platforms (C.1, P.3); and b) maintaining ultra reliability during UAV trajectory changes (C.4) and remote AIoT device operations (P.2). To overcome these issues, the authors developed a low complexity optimization framework that jointly manages resource allocation and UAV trajectory planning. The optimization is based on a successive convex approximation algorithm combined with a random restart method to escape local minima and improve convergence. The model aims to minimize the weighted sum of energy consumption across UAVs and ground AIoT devices while satisfying URLLC level constraints, such as sub-millisecond latency and ultra low packet error rates. Specifically, smaller blocklengths and tighter error tolerances are incorporated into the QoS formulation to ensure deterministic transmission bounds. The proposed work has been shown to reduces total energy consumption (P.3) compared to benchmark fixed trajectory and static resource allocation strategies. This work enhances URLLC support in dynamic edge environments by ensuring end-to-end reliability (P.2), achieving latency constrained energy savings (P.3). These capabilities make the proposed framework suitable for real-time Industry 5.0 deployments involving mobile AIoT networks and mission critical edge applications.
@@
In [63], Imtiaz et al. proposed a DT enabled MEC framework aimed at improving computational energy efficiency in latency sensitive consumer Industrial IoT (IIoT) applications for Industry 5.0 smart manufacturing environments (A.5) to balance latency vs. reliability trade-offs (C.1). The system as shown in Figure 7 introduces predictive digital twin agents that continuously synchronize with physical edge computing nodes, enabling proactive decision making for task offloading. Unlike traditional reactive MEC systems, the proposed framework predicts resource bottlenecks in advance and dynamically schedules tasks to minimize latency and energy consumption simultaneously (T.5, T.1).
@@
The key optimization challenge is modeled as a mixed integer non-linear programming problem that jointly minimizes energy consumption and task latency while satisfying strict URLLC constraints. To achieve computational tractability, the authors apply piecewise linear approximation to relax non-linear constraints. Task assignment decisions are made based on predicted CPU availability, wireless channel quality, and residual energy of nodes, allowing latency aware offloading while ensuring reliability. This predictive scheduling framework enhances both the responsiveness and sustainability of real-time IIoT applications by allocating edge resources more efficiently (P.3). The proposed work has been shown to improve computational energy efficiency compared to traditional edge only models and reduces the service outage probability under strict latency and reliability requirements. These improvements contribute to end-to-end latency reduction (P.1), reliability improvement (P.2), and resource optimization (P.3) which are, key performance enhancements required in URLLC systems. The proposed DT assisted optimization approach provides energy efficient foundation for deploying URLLC enabled smart manufacturing services in Industry 5.0.
@@
Li et al. [55] proposed a low latency terminal mutual access architecture based on 5G LAN cross-domain networking designed to support real-time communication demands in dynamic industrial scenarios. The architecture addresses the shortcomings of traditional communication frameworks that route data through central servers, introducing latency and single points of failure. Instead, their solution enables D2D communication using 5G LAN capabilities, which is important for latency sensitive applications such as remote surgeries (A.3), and supporting the evolution toward next-generation mobile networks (A.6)
The proposed architecture addresses two main challenges: a) enabling real-time, ultra reliable communication across multiple 5G LAN domains managed by different operators (C.2); and b) maintaining privacy and security while bypassing centralized cores (C.3). The architecture integrates MEC (T.1) to host computation intensive tasks near end users, network slicing (T.2) to logically separate services, and NFV and SFC (T.3) to flexibly deploy and interconnect virtual network functions across domains. To achieve this, the authors use a distributed mesh topology with user plane function based routing (i.e., decentralized traffic anchoring at the edge to enable direct data forwarding between terminals without passing through the core network) supported by secure access service edge and edge offloading (T.1). These mechanisms enable traffic to bypass base stations and core networks (C.4), thereby achieving latency isolation and privacy preserving communication under a multi-domain setup (C.3).
@@ IV-F Cross-Layer Networking Solutions
Gao et al. [68] addressed the challenge of delivering URLLC compliant services for latency sensitive metaverse applications such as real-time holographic communication (A.4). The primary difficulty lies in simultaneously satisfying ultra-low latency, high reliability, and immersive QoE requirements under limited wireless and computing resources (C.1, C.4). To address this, the authors proposed a multi-objective optimization framework for joint allocation of communication, computation, and storage resources in edge assisted networks (T.1). The system model considers a heterogeneous user base with varying latency budgets, traffic demands, and task data sizes. Reliability is enforced using probabilistic bounds on decoding errors  (P.1, P.2). The resulting non-convex, integer constrained problem is solved using a decomposition based evolutionary algorithm to approximate the Pareto front, enabling designers to select operating points that best balance URLLC constraints and QoE. Simulation results demonstrate improved delay guarantees (P.1) and higher adaptability (P.3).
@@
In dynamic next-generation wireless networks (A.6), accurate channel prediction is fundamental for optimizing communication and ensuring low latency, reliability and security of data transmission. AI, Including DL and RL, offer great advantages for URLLC use cases. For instance, Li et al. [54] proposed a DRL based joint scheduler that dynamically balances QoS trade-offs between eMBB and URLLC (A.6). Unlike static or risk sensitive models, the authors formulated a multi-objective optimization problem using reliability (P.2) and service satisfaction level as key metrics. Reliability is quantified through the probability of successful URLLC packet decoding under finite blocklength constraints, where decoding error probabilities exceeding a target threshold (e.g., $10^{-5}$) adds penalties in the reward. The environment is modeled such that the state includes channel quality indicators and queue lengths, the action consists of selecting resource block assignments and puncturing decisions for URLLC packets, and the reward function jointly optimizes service satisfaction level fairness for eMBB and packet error rate constraints for URLLC traffic. They used Deep Deterministic Policy Gradient (DDPG) with prioritized experience replay to train the scheduler on real-time traffic variations. The proposed approach has been shown to increase service satisfaction level (P.3) and maintain reliability requirements for URLLC transmissions (P.2). By using DRL (T.4), the scheduler learns to reallocate resources under stochastic traffic demands and fluctuating wireless channel conditions (C.4), thereby enabling latency aware, priority preserving access in mixed traffic scenarios. This supports the goals of URLLC by ensuring intelligent resource orchestration, reduced queuing delays, and improved service continuity under hard latency and reliability constraints (P.1).
@@
Similar to [54] and [70], She et al. [34] proposed a multi-level DL architecture that addresses the challenge of meeting URLLC performance requirements (C.4) in non-stationary 6G networks (A.6). Figure 11 shows the proposed framework, which integrates device level intelligence for mobility prediction, edge level (T.1) intelligence for scheduler design, and cloud level intelligence for user association. At the device level, they use fully connected DNNs trained on tactile mobility traces to predict user movements, achieving a prediction error probability below $10^{-5}$, thereby enhancing reliability guarantees (P.2). At the cell level, a DRL based actor critic scheduler is developed and evaluated in a live LTE testbed. The scheduler is trained using two DNNs (i.e., actor and critic), each with multiple hidden layers, to minimize packet loss and ensure delay and jitter bounds. They demonstrated real-time performance under strict constraints using pre-training in a digital twin, then fine tuning in real-world deployments (T.5, P.1). At the network level, the system uses DNN based user association, trained with large scale channel gains and packet arrival rates, and compared against optimal and baseline models. To address learning in dynamic environments, the authors employed deep transfer learning to adapt pre-trained models with limited new data (T.4), and applied hierarchical federated learning to enable distributed training while preserving privacy (C.3) and reducing communication overhead (C.4). The proposed approach has been shown to improve resource utilization (P.3) and energy efficient user association across varying service types and user distributions (P.4, P.5). By extending the learning from device to edge, and cloud layers, this study supports the main principles of URLLC by enabling ultra low latency task scheduling, accurate user mobility tracking, and proactive resource orchestration in highly dynamic environments which ensures reliable performance across heterogeneous traffic scenarios (C.1).
@@
Zhou et al. [28] addressed the challenge of supporting URLLC for smart healthcare services in Internet of Health Things (IoHT) networks (A.3, A.6). Ensuring ultra low latency and high reliability for tasks such as remote health monitoring and emergency diagnosis is challenging due to device heterogeneity, dynamic channel conditions, constrained computation (C.1, C.4), and the need for privacy preserving solutions (C.3). The authors proposed a learning based task offloading approach that uses federated DRL model to optimize computation offloading decisions under URLLC constraints. The model is based on the exponential weight algorithm for exploration and exploitation to optimize computation offloading decisions under URLLC constraints. The proposed approach adopts a hybrid architecture where IoHT devices decide whether to locally execute or offload tasks to edge servers. A federated DQN is used to train local agents collaboratively without sharing raw data, preserving patient privacy (C.3). Each agent observes a state composed of local queue length, channel quality, and device energy. The action includes binary offloading decisions and edge server selections. The reward function includes task success rate and delay violations in order to enforce URLLC compliance. Agents update local policies using prioritized experience replay and synchronize with a global model to adapt to environmental changes (T.4). The proposed approach has been shown to reduce average task latency (P.1) and improve offloading success rate (P.2) compared to traditional approaches. This enables URLLC support for delay sensitive healthcare applications such as real-time ECG monitoring and AR assisted surgery. By integrating privacy aware learning, energy-efficient offloading, and delay sensitive scheduling, the work contributes to reliable and low latency healthcare delivery in dense IoHT environments.
@@
Kokkinis et al. [49] addressed the challenge of synchronizing video and haptic streams under URLLC constraints in tactile Internet environments, which falls under the application domain of Remote Control and Haptics (A.4). A key issue they focused on was meeting distinct service level requirements for haptic and video traffic, such as sub-10 ms latency for haptic control loops and desynchronization thresholds below 50 ms for user perception, while maintaining spectral efficiency and robustness amidst fluctuating wireless conditions. These challenges inherently involve latency vs. reliability trade-offs (C.1) and mobility management (C.4). The authors proposed a soft actor-critic based Deep Reinforcement Learning (DRL) framework, leveraging AI/ML for URLLC (T.4), to optimize radio resource allocation between haptic (URLLC slice) and video (eMBB slice) modalities. The DRL environment models state by capturing average latency, packet loss, data rate, spectrum efficiency, and buffer occupancy per slice, with actions corresponding to discrete allocations of resource blocks. Their reward function penalizes violations of service-specific constraints, including slice-wise data rate (P.3), packet loss (P.2), and inter-modal synchronization (P.1). The system reduces computational complexity in edge deployments by performing inter-slice scheduling every 50 ms instead of 1 ms. The proposed framework has been shown to achieve a higher satisfaction rate (P.4) under high user density (P.5) and spectrum efficiency fluctuations compared to traditional approaches. By modeling heterogeneous service demands and optimizing resource allocation via DRL, their approach enables stable, low latency, and synchronized teleoperation services, contributing to the evolving capabilities of next-generation mobile networks (A.6).
@@
The above studies show that AI and ML techniques are increasingly central to enabling URLLC in 6G enabled Industry 5.0 applications. From optimizing dynamic slicing and uplink scheduling to managing delay constrained offloading and haptic video synchronization, AI agents provide scalable and adaptive control policies under stochastic traffic and wireless dynamics (T.4). These learning based approaches: a) improve latency reliability trade-offs (C.1); and b) supports intelligent orchestration across mobility intensive (C.4) and heterogeneous service environments. This is done by bridging real-time perception, decision-making, and control, in order to contribute to fulfilling the end-to-end URLLC objectives of latency reduction (P.1), reliability improvement (P.2), and resource efficiency (P.3, P.4), which in turn makes it one of the main enablers of intelligent and responsive Industry 5.0 systems.